Rochester Heating Systems, Inc. calculates its predetermined overhead rate on a quarterly basis. The following estimates were made for the current year Quarterly Predetermined Overhead Estimated Estimated Rate (per Manufacturing Direct-Labor direct-labor Hours 25,000 16,000 Overhead hour) First quarter 75,000 Second 96,000 quarter Third quarter 37,500 Fourth 12,500 84,000 14,000 quarter Total $292,500 67,500 The firms main product, part number SC71, requires $200 of direct material and 20 hours of direct labor per unit. The labor rate is $15 per hour
Answer #1

Part 1 Calculation of predetermined overhead rate:

Overhead rate = estimated overhead / direct labour hours.

Quarter 1 rate = 75000 / 25000 = 3 (same for quarter 3)

Quarter 2 rate = 96000 / 16000 = 6 (same for quarter 4)

Part 2 Calculation of price per unit of SC71:

Particulars Direct Material (A) Direct labour (B) Quarter 1 (February) Quarter 2 (Ma 200 300 200 300 (20X 15) (20 X 15) Total (A B) Manufacturing overhead 500 60 500 120 (20 X 3) (20 X 6) Total cost Marku Price of SC71 560 56 616 620 62 68

Therefore price for the months of February and May are 616 and 682 respectively